# MCMC Marginalisation Bias and $\Lambda$CDM tensions

### v2

@@ 4 An MCMC puzzle
It should be evident that a degeneracy (banana-shaped contour) in the 2D MCMC posterior, which gives rise to a projection effect in 1D posteriors, explains differences between posterior peaks and best fits in Table 1. Moreover, the effect of this degeneracy starts to impact results at relatively low redshifts, i. e. $z_{\textrm{min}}=0.7$. More puzzling still, as is clear from Fig. 6, by $z_{\textrm{min}}=1.2$ any preference in the data for a higher $H_{0}$ value has been lost. All that remains is a bump in the $H_{0}$ posterior that coincides with the best fit, $H_{0}=154.35$ km/s/Mpc (dashed line). Nevertheless, returning to the raw data, it is blatantly clear that one can fit an almost horizontal line to CC data with $z>1.2$. Moreover, one can get such a function from the $\Lambda$CDM model (1) through a small value of $\Omega_{m}$ and a large value of $H_{0}$, and there is no reason to believe that the data poorly constrains the model. Nevertheless, this is the conclusion a cosmologist reaches if they rely exclusively on MCMC for their ground truth.
@@ 5 A tension with Planck
We again marginalise over the free parameters $(H_{0},\Omega_{m},r_{d})$ with MCMC. In Fig. 8 we present the posteriors. While $r_{d}$ is Gaussian and peaked on our Planck prior, as expected, the $\Omega_{m}$ posterior is peaked at $\Omega_{m}\sim 0.6$ and the fact that the fall off in the distribution is gradual beyond the peak leads to a pile up of configurations in the top left corner of the $(H_{0},\Omega_{m})$-plane. This fall off continues beyond $\Omega_{m}=1$ and if the prior is relaxed, the $H_{0}$ peak shifts to smaller values. So, once again all the hallmarks of projection effects are present. That being said, given the sharp fall off in the $\Omega_{m}$ distribution to smaller $\Omega_{m}$ values, some discrepancy appears to be evident with the Planck values (dashed lines).
@@
We now run the MCMC chain through our PD methodology. From Fig. 9, we can see that the $R(H_{0})$ and $R(\Omega_{m})$ distributions prefer smaller values of $H_{0}$ and larger values of $\Omega_{m}$. The peak of the distributions occurs at $H_{0}=42.40$ km/s/Mpc and $\Omega_{m}=0.795$. The lone dot in the $R(H_{0})$ distribution at low values of $H_{0}$ tells us that the distribution falls off sharply below $H_{0}=40$ km/s/Mpc. Note, since we employed generous uniform priors $H_{0}\in[0,200]$, the priors are not impacting the $R(H_{0})$ distribution, so it is expected that the distribution falls off to zero on both sides. In contrast, the $R(\Omega_{m})$ distribution is one-sided and fails to fall off in the direction of larger values within the uniform priors $\Omega_{m}\in[0,1]$. The tension with Planck falls between $95\%$ ($2\sigma$) and $99.7\%$ ($3\sigma$) confidence intervals. By integrating the PDF as far as the black lines corresponding to the Planck values in Fig. 9, we estimate that the Planck $H_{0}$ is located at the $96\%$ ($2.1\sigma$) confidence level, while the Planck $\Omega_{m}$ value is at the $98.7\%$ ($2.5\sigma$) confidence level.
@@ 6 Concluding remarks
A $\chi^{2}$ likelihood is a metric or measure of how well a model fits data. The point in model parameter space that fits the data the best possesses the lowest $\chi^{2}$. Once one has identified this point, the problem remains to establish $68\%$, $95\%$, etc, confidence intervals in parameter space. In cosmology and astrophysics, MCMC is the prevailing technique for estimating confidence intervals. Its great advantage is that it allows one to i) globally sample the parameter space and ii) arrive at posteriors that serve as an estimate of the errors even with non-Gaussian distributions. In contrast, if one minimises the $\chi^{2}$ by gradient descent, there is always a risk that one ends up in a local minimum, i. e. the global minimum is missed, while error estimation through Fisher matrix assumes any distribution is Gaussian. The appeal of MCMC marginalisation is that it is widely applicable. However, the point of this paper is that limitations exist, even in the simplest model.
@@
Indeed, what happens when the MCMC posterior is impacted by a degeneracy and no longer tracks points in parameter space that fit the data the best? Traditionally, volume effects are seen as the preserve of higher-dimensional models, e. g. Herold:2021ksg ; Gomez-Valent:2022hkb ; Meiers:2023gft , but projection effects also occur in the minimal $\Lambda$CDM model when one fits the model to data binned by redshift in the late Universe Colgain:2022tql . As explained in Colgain:2022tql , this “projection effect” is driven by OHD, $H(z_{i})$, and angular diameter or luminosity distance data, $D_{A}(z_{i})$ or $D_{L}(z_{i})$, respectively only constraining the combinations $\Omega_{m}h^{2}$ and $(1-\Omega_{m})h^{2}$ well, with high redshift data $z_{i}\gg 0$. In practice, this restricts MCMC configurations to constant $\Omega_{m}h^{2}$ and constant $(1-\Omega_{m})h^{2}$ curves in the $(H_{0},\Omega_{m})$ plane, and as the curves stretch due to DE or matter being less well constrained in high redshift bins, projection effects lead to shifts in the peaks of MCMC posteriors and the emergence of non-Gaussian tails Colgain:2022tql . We stress that one sees the same effect in PDFs of best fit $(H_{0},\Omega_{m})$ parameters in a large number of mock data realisations Colgain:2022tql , so the problem is more general than MCMC; there is an inherent bias in the $\Lambda$CDM model when one fits it to redshift binned $H(z)$ or $D_{A}(z)$ or $D_{L}(z)$ data. Within MCMC, one sees this effect in the errors, the shift in peaks in 1D posteriors, but also in the drift of the parameters corresponding to the $\chi^{2}$ minimum outside of the $68\%$ confidence intervals. Highlighting this (expected) bias in MCMC using OHD is the opening salvo (result) of this paper.
@@
Why should one care? This is evidently only a problem if one bins data and confronts the $\Lambda$CDM model. First, note that some data sets are inherently binned. For example, effective redshifts are assigned to CC and BAO analysed in a given redshift bin, while each strongly lensed system constitutes its own bin. Working with binned data is unavoidable. Secondly, $\Lambda$CDM tensions point to a problem with the $\Lambda$CDM model once the tensions become widespread and persistent. As explained in Krishnan:2020vaf , if the minimal $\Lambda$CDM model is too simple, one expects redshift evolution of $\Lambda$CDM cosmological parameters as it is confronted to redshift binned data. Hints of these trends are now evident in $H_{0}$ Wong:2019kwg ; Millon:2019slk ; Dainotti:2021pqg ; Colgain:2022nlb ; Colgain:2022rxy ; Malekjani:2023dky ; Hu:2022kes ; Jia:2022ycc ; Krishnan:2020obg ; Dainotti:2022bzg , $\Omega_{m}$ Risaliti:2015zla ; Risaliti:2018reu ; Lusso:2020pdb ; Yang:2019vgk ; Khadka:2020vlh ; Khadka:2020tlm ; Khadka:2021xcc ; Pourojaghi:2022zrh ; Colgain:2022nlb ; Colgain:2022rxy ; Malekjani:2023dky ; Pasten:2023rpc ; Sakr:2023hrl and $S_{8}$/$\sigma_{8}$ Esposito:2022plo ; Adil:2023jtu ; ACT:2023dou ; ACT:2023kun (also Miyatake:2021qjr ; Alonso:2023guh ) across a host of different observables. This evolution is an expected hallmark of model breakdown, which must happen at some redshift if systematics are not universally at play.
@@
The main problem with redshift dependent $\Lambda$CDM cosmological parameters  [^11] There is a separate interpretation problem as the cosmology literature works with parameters “defined today”. In more mathematical language, this is simply the statement that one solves an ordinary differential equation (ODE), namely the Friedmann equation or equivalent, by specifying an integration constant, e.g. $H_{0}=H(z=0)$ or $\rho_{m}(z=0)=\rho_{m0}=H_{0}^{2}\Omega_{m}$. However, this is a mathematical statement and it still needs to be confirmed observationally that $H_{0}$ or $\rho_{m0}$ are bona fide constants. This cannot be a priori assumed, because it is mathematical prediction of the model. If the model is correct, a constant $H_{0}$ and $\Omega_{m}$ will be supported by the data. See Krishnan:2020vaf for further discussion. is one needs to assign a statistical significance to any trend. At a purely practical level, this entails constructing bins centered on different redshifts and identifying discrepancies in $\Lambda$CDM parameters between bins, ideally in the same observable, so that the potential systematics are under greatest control. As demonstrated both mathematically and observationally with the CC data in section 2, MCMC marginalisation leads to biased inferences when one bins the data. In this paper we have resorted to profile distributions Gomez-Valent:2022hkb to overcome this bias and have applied the technique to a setting where projected 1D $\Lambda$CDM posteriors are expected to be non-Gaussian for the reasons outlined above and in section 2. This new technique, provides a complementary perspective that confirms the least square fits of observed and mock data presented in Colgain:2022nlb ; Colgain:2022rxy ; Malekjani:2023dky , where evidence for redshift evolution in $H_{0}$ and $\Omega_{m}$ was presented. Regardless of the methodology, the objective is to drill down on the prevailing assumption that cosmological parameters are constants. In the era of tensions in cosmology, nothing can be assumed, especially noting that the tensions are in essence showing an example of evolution of these parameters with redshift and hence the model breakdown.
@@
More concretely, in this paper with both mock simulations and profile distributions we have shown that high redshift CC data has a preference for a non-evolving $H(z)$ over Planck-$\Lambda$CDM at approximately $\sim 2\sigma$. This trend, which constitutes the second result of the paper, is unquestionable, as it is visible in the data. Note, we have not propagated systematic uncertainties, so the significance will be less when these are properly propagate. Nevertheless, low and high redshift CC data currently have a preference for different $\Lambda$CDM cosmological parameters. Interestingly, as we have shown in section 4, resorting exclusively to MCMC one cannot arrive at this conclusion because 2D posteriors are unconstrained. It is important that observables at all redshifts return similar $H_{0}$ values because if the CC program is claiming an 8% constraint on the Hubble constant, $H_{0}=66.7\pm 5.5$ km/s/Mpc Moresco:2023zys , it is imperative that all subsets of the data are consistent with this result. If they are not, then we are staring at either systematics or model breakdown. Admittedly, demanding self-consistency of subsets of a data set confronted to a model is a high bar, but it is important that data sets result in overlapping constraints on $\Lambda$CDM parameters, otherwise this makes cosmological inferences moot. Note, the $\Lambda$CDM model is largely only well tested in the DE dominated regime $z\lesssim 1$ and at very high redshifts $z\sim 1100$, which leaves a wide expanse of redshifts to be explored in order to confirm or refute the model. Given the existing $\Lambda$CDM tensions Perivolaropoulos:2021jda ; Abdalla:2022yfr , and the hints of evolution in $H_{0}$, $\Omega_{m}$ and $S_{8}$ across assorted probes in the late Universe $z\lesssim 5$, it would be surprising if all discrepancies could be explained away by systematics. [^12] We are open to the possibility, we just consider it a bad bet at the moment. The odds can of course change as observations improve.
@@
As an aside, it is intriguing that CC data has a preference for larger $H_{0}$ best fit values and smaller $\Omega_{m}$ best fit values beyond $z_{\textrm{min}}=0.7$, as this is traditionally the transition redshift between decelerated and accelerated expansion. Moreover, at higher redshifts $z\sim 2.3$, there is not only a longstanding anomaly in Lyman-$\alpha$ BAO duMasdesBourboux:2020pck , but QSOs also show a preference for a lower luminosity distance, $D_{L}(z)$, relative to Planck-$\Lambda$CDM Risaliti:2015zla ; Risaliti:2018reu . Translated into $\Lambda$CDM parameters, this corresponds to conversely larger $\Omega_{m}$ values, e. g. Yang:2019vgk ; Khadka:2020vlh ; Khadka:2020tlm ; Khadka:2021xcc ; Pourojaghi:2022zrh , and consequently smaller $H_{0}$ values. Thus, the emerging probes CC and QSOs Moresco:2022phi do not appear to be in sync on high redshift $\Lambda$CDM inferences. Nevertheless, neither may be inconsistent with the anomaly in Lyman-$\alpha$ BAO. Relative to Planck-$\Lambda$CDM, Lyman-$\alpha$ BAO prefers smaller values of $D_{M}(z):=c\int_{0}^{z}1/H(z^{\prime})\,\textrm{d}z$ and smaller values of $H(z)$ (larger values of $D_{H}(z):=c/H(z)$). [^13] In this statement we assumed the Planck value $r_{d}\sim 147$ Mpc Planck:2018vyg If we reinstate the radius of the sound horizon in these expressions, one recognises that changing the sound horizon, as advocated by early Universe resolutions to Hubble tension, cannot consistently address the Lyman-$\alpha$ BAO anomaly. In general, even for the Planck-$\Lambda$CDM sound horizon, one cannot get both a smaller $D_{M}(z)$ and smaller $H(z)$ from a strictly increasing function, such as the $\Lambda$CDM $H(z)$. As a result, deviations from the Planck-$\Lambda$CDM model that address this anomaly are expected to lead to wiggles in $H(z)$ Akarsu:2022lhx , which are unsurprisingly seen in data reconstructions Zhao:2017cud ; Wang:2018fng ; Escamilla:2021uoj . Finally, evolution in $H_{0},\Omega_{m}$ discussed here cannot be explained or accommodated by early resolutions to Hubble tension relying on a change in the $r_{d}$ at very high $z$.. If CC data prefer less evolution in $H(z)$ in the matter-dominated regime, then this is consistent with the preference for a smaller $H(z)$ from Lyman-$\alpha$ BAO. Furthermore, QSO data prefers smaller luminosity distances $D_{L}(z)$ relative to Planck, which are consistent with the smaller $D_{M}(z)\propto D_{L}(z)$ values preferred by Lyman-$\alpha$ BAO. Thus, even if CC and QSOs appear to be showing diverging behaviour in the cosmological parameters $(H_{0},\Omega_{m})$, this may still turn out to be consistent with Lyman-$\alpha$ BAO. We await future DESI DESI:2023ytc data releases to ascertain if the non-evolving $H(z)$ trend in high redshift CC data is physical or not.
@@
Finally, we come to our third and main result outlined in section 5. We have revisited a $\sim 2\sigma$ tension between high redshift CC and BAO data reported in Colgain:2022rxy , where the significance was estimated through mock simulations. Here, we have upgraded the BAO data to the latest constraints and again recover a $>2\sigma$ discrepancy in $(H_{0},\Omega_{m})$ with different methodology. This provides a consistency check that there is evolution in OHD between low and high redshifts in the late Universe. Note, this evolution runs contrary to the non-evolving $H(z)$ seen in high redshift CC data because it assumes Planck has accurately constrained the high redshift behaviour of the Hubble parameter in (1). Nevertheless, both with and without a Planck prior on $\Omega_{m}h^{2}$, evolution at $\gtrsim 2\sigma$ is evident in OHD data. It should be stressed that evolution is evident in PDFs of best fit $\Lambda$CDM parameters fitted to a large number of Planck-$\Lambda$CDM mocks Colgain:2022tql , so evolution in observed data can be expected. It is imperative to revisit the remaining observations in Colgain:2022rxy ; Malekjani:2023dky in order to confirm the significance of $\sim 2\sigma$ hints of evolution found separately in Type Ia SN and QSO data sets.

### v3

@@ 4 An MCMC puzzle
It should be evident that a degeneracy (banana-shaped contour) in the 2D MCMC posterior, which gives rise to a projection effect in 1D posteriors, explains differences between posterior peaks and best fits in Table 1. Moreover, the effect of this degeneracy starts to impact results at relatively low redshifts, i. e. $z_{\textrm{min}}=0.7$. More puzzling still, as is clear from Fig. 7, by $z_{\textrm{min}}=1.2$ any preference in the data for a higher $H_{0}$ value has been lost  [^10] We use emcee on its default algorithm. It is plausible that changing the algorithm has a bearing on the outcome, but it is not expected to remove the degeneracy, since the $\Lambda$CDM model must transition to an effective 1D model in high redshift bins. This underscores Bayesian cosmology intuition that high redshift data only constrains the combination $\Omega_{m}h^{2}$. The point here is that intuition rests upon experience with MCMC, so Fig. 7 is expected regardless of the algorithm.. All that remains is a bump in the $H_{0}$ posterior that coincides with the best fit, $H_{0}=154.35$ km/s/Mpc (dashed line). Nevertheless, returning to the raw data, it is blatantly clear that one can fit an almost horizontal line to CC data with $z>1.2$. Moreover, one can get such a function from the $\Lambda$CDM model (1) through a small value of $\Omega_{m}$ and a large value of $H_{0}$, and there is no reason to believe that the data poorly constrains the model. Nevertheless, this is the conclusion a cosmologist reaches if they rely exclusively on MCMC marginalisation for their ground truth. Note, we do not need to construct a profile likelihood or distribution to arrive at this conclusion; it suffices to visibly inspect the data and understand that the $\Lambda$CDM model reduces to $H(z)=\textrm{constant}$ in the limit $\Omega_{m}\rightarrow 0$.
@@ 5 A tension with Planck
We again marginalise over the free parameters $(H_{0},\Omega_{m},r_{d})$ with MCMC. In Fig. 9 we present the posteriors. While $r_{d}$ is Gaussian and peaked on our Planck prior, as expected, the $\Omega_{m}$ posterior is peaked at $\Omega_{m}\sim 0.6$ and the fact that the fall off in the distribution is gradual beyond the peak leads to a pile up of configurations in the top left corner of the $(H_{0},\Omega_{m})$-plane. This fall off continues beyond $\Omega_{m}=1$ and if the prior is relaxed, the $H_{0}$ peak shifts to smaller values. So, once again all the hallmarks of projection effects are present. That being said, given the sharp fall off in the $\Omega_{m}$ distribution to smaller $\Omega_{m}$ values, some discrepancy appears to be evident with the Planck values (dashed lines). We remind the reader again that the $H_{0}$ posterior is unimpacted by our prior $H_{0}\in[0,200]$. However, we note that the 2D posterior in the $(H_{0},\Omega_{m})$-plane is largely unconstrained, so one suspects that projection effects are present in the $H_{0}$ posterior. What distinguishes the corner plot in Fig. 9 from the corner plots in Fig.2 and Fig. 7 as genuine is that a peak in the $H_{0}$ posterior at smaller values coincides with a peak in the $\Omega_{m}$ posterior at larger values, thereby recovering the expected $H_{0}-\Omega_{m}$ anti-correlation in the $\Lambda$CDM model. In Fig. 2, where the focus is on the secondary phantom peak, and Fig. 7 this anti-correlation is not evident in the $H_{0}$ and $\Omega_{m}$ posteriors.
@@
We now run the MCMC chain through our PD methodology. From Fig. 10, we can see that the $R(H_{0})$ and $R(\Omega_{m})$ distributions prefer smaller values of $H_{0}$ and larger values of $\Omega_{m}$. The peak of the distributions occurs at $H_{0}=42.40$ km/s/Mpc and $\Omega_{m}=0.795$. The lone dot in the $R(H_{0})$ distribution at low values of $H_{0}$ tells us that the distribution falls off sharply below $H_{0}=40$ km/s/Mpc. Note, since we employed generous uniform priors $H_{0}\in[0,200]$, the priors are not impacting the $R(H_{0})$ distribution, so it is expected that the distribution falls off to zero on both sides. In contrast, the $R(\Omega_{m})$ distribution is one-sided and fails to fall off in the direction of larger values within the uniform priors $\Omega_{m}\in[0,1]$. The tension with Planck falls between $95\%$ ($2\sigma$) and $99.7\%$ ($3\sigma$) confidence intervals. By integrating the PDF as far as the black lines corresponding to the Planck values in Fig. 10, we estimate that the Planck $H_{0}$ is located at the $96\%$ ($2.1\sigma$) confidence level, while the Planck $\Omega_{m}$ value is at the $98.7\%$ ($2.5\sigma$) confidence level. We remind the reader that some care is required with $\Omega_{m}$ inferences because the PD is curtailed by the bounds. The same is true for MCMC posteriors with priors.
@@ 6 Concluding remarks
A $\chi^{2}$ likelihood is a metric or measure of how well a model fits data. The point in model parameter space that fits the data the best possesses the lowest $\chi^{2}$. Once one has identified this point, the problem remains to establish $68\%$, $95\%$, etc, confidence intervals in parameter space. In cosmology and astrophysics, MCMC is the prevailing technique for estimating credible intervals. Its great advantage is that it allows one to i) globally sample the parameter space and ii) arrive at posteriors that serve as an estimate of the errors even with non-Gaussian distributions. In contrast, if one minimises the $\chi^{2}$ by gradient descent, there is always a risk that one ends up in a local minimum, i. e. the global minimum is missed, while error estimation through Fisher matrix essentially assumes any distribution is Gaussian. The appeal of MCMC marginalisation is that it is widely applicable. However, the point of this paper is that limitations exist, even in the simplest model.
@@
Indeed, what happens when the MCMC posterior is impacted by a degeneracy and no longer tracks points in parameter space that fit the data the best? Traditionally, volume effects are seen as the preserve of higher-dimensional models, e. g. Herold:2021ksg ; Gomez-Valent:2022hkb ; Meiers:2023gft , but projection effects also occur in the minimal $\Lambda$CDM model when one fits the model to data binned by redshift in the late Universe Colgain:2022tql . As explained in Colgain:2022tql , this “projection effect” is driven by OHD, $H(z_{i})$, and angular diameter or luminosity distance data, $D_{A}(z_{i})$ or $D_{L}(z_{i})$, respectively only constraining the combinations $\Omega_{m}h^{2}$ and $(1-\Omega_{m})h^{2}$ well, with high redshift data $z_{i}\gg 0$. In practice, this restricts MCMC configurations to constant $\Omega_{m}h^{2}$ and constant $(1-\Omega_{m})h^{2}$ curves in the $(H_{0},\Omega_{m})$ plane, and as the curves stretch due to DE or matter being less well constrained in high redshift bins, projection effects lead to shifts in the peaks of MCMC posteriors and the emergence of non-Gaussian tails Colgain:2022tql . We stress that one sees the same effect in PDFs of best fit $(H_{0},\Omega_{m})$ parameters in a large number of mock data realisations Colgain:2022tql , so the problem is more general than MCMC marginalisation; there is an inherent bias in the $\Lambda$CDM model when one fits it to redshift binned $H(z)$ or $D_{A}(z)$ or $D_{L}(z)$ data. Succinctly put, the $\Lambda$CDM model confronted to redshift-binned OHD must transition from a 2D model to an effective 1D model. This impacts both Bayesian and frequentist methods, but the effect is more pronounced in MCMC marginalisation. Within MCMC marginalisation, one sees this effect in the errors, the shift in peaks in 1D posteriors, but also in the drift of the parameters corresponding to the $\chi^{2}$ minimum outside of the $68\%$ credible intervals. Highlighting this (expected) feature in MCMC marginalisation using OHD is the opening salvo (result) of this paper.
@@
Why should one care? This is evidently only a problem if one bins data and confronts the $\Lambda$CDM model. First, note that some data sets are inherently binned. For example, effective redshifts are assigned to CC and BAO analysed in a given redshift bin, while each strongly lensed system constitutes its own bin. Working with binned data is unavoidable. Secondly, $\Lambda$CDM tensions point to a problem with the $\Lambda$CDM model once the tensions become widespread and persistent. As explained in Krishnan:2020vaf , if the minimal $\Lambda$CDM model is too simple, one expects redshift evolution of $\Lambda$CDM fitting parameters as it is confronted to redshift binned data. Hints of these trends are now evident in $H_{0}$ Wong:2019kwg ; Millon:2019slk ; Dainotti:2021pqg ; Colgain:2022nlb ; Colgain:2022rxy ; Malekjani:2023dky ; Hu:2022kes ; Jia:2022ycc ; Krishnan:2020obg ; Dainotti:2022bzg , $\Omega_{m}$ Risaliti:2015zla ; Risaliti:2018reu ; Lusso:2020pdb ; Yang:2019vgk ; Khadka:2020vlh ; Khadka:2020tlm ; Khadka:2021xcc ; Pourojaghi:2022zrh ; Colgain:2022nlb ; Colgain:2022rxy ; Malekjani:2023dky ; Pasten:2023rpc ; Sakr:2023hrl and $S_{8}$/$\sigma_{8}$ Esposito:2022plo ; Adil:2023jtu ; ACT:2023dou ; ACT:2023kun (also Miyatake:2021qjr ; Alonso:2023guh ) across a host of different observables. This evolution in fitting parameters is an expected hallmark of model breakdown, which must happen at some redshift if systematics are not universally at play.
@@
The main problem with redshift dependent $\Lambda$CDM fitting parameters  [^12] There is a separate interpretation problem as the cosmology literature works with parameters “defined today”. In more mathematical language, as explained in the introduction, this is simply the statement that one solves an ordinary differential equation (ODE), namely the Friedmann equation or equivalent, by specifying an integration constant, e.g. $H_{0}=H(z=0)$ or $\rho_{m}(z=0)=\rho_{m0}=H_{0}^{2}\Omega_{m}$. However, this is a mathematical statement and it still needs to be confirmed observationally that $H_{0}$ or $\rho_{m0}$ are bona fide constants from the perspective of observation. This cannot be a priori assumed, because it is mathematical prediction of the model. If the model is correct, confronting the $H_{0}$ and $\Omega_{m}$ fitting parameters to data at different redshifts will return consistent constant values of the parameters. This is an important observational consistency test of the $\Lambda$CDM model. See Krishnan:2020vaf for further discussion. is one needs to assign a statistical significance to any trend. At a purely practical level, this entails constructing bins centered on different redshifts and identifying discrepancies in $\Lambda$CDM fitting parameters between bins, ideally in the same observable, so that the potential systematics are under greatest control. As demonstrated both mathematically and observationally with the CC data in section 2, MCMC marginalisation leads to biased inferences from projection effects when one bins the data. In this paper we have resorted to profile distributions Gomez-Valent:2022hkb to overcome this bias and have applied the technique to a setting where projected 1D $\Lambda$CDM posteriors are expected to be non-Gaussian for the reasons outlined above and in section 2. This new technique, provides a complementary perspective that confirms the least square fits of observed and mock data presented in Colgain:2022nlb ; Colgain:2022rxy ; Malekjani:2023dky , where evidence for redshift evolution in $H_{0}$ and $\Omega_{m}$ was presented. Regardless of the methodology, the objective is to drill down on the prevailing assumption that $\Lambda$CDM fitting parameters are constants. In the era of tensions in cosmology, nothing can be assumed, especially noting that the tensions are in essence showing an example of evolution of these fitting parameters with redshift and hence the model breakdown.
@@
More concretely, in this paper with both mock simulations and profile distributions we have shown that high redshift CC data has a preference for a non-evolving $H(z)$ over Planck-$\Lambda$CDM at approximately $\sim 2\sigma$. This trend, which constitutes the second result of the paper, is unquestionable, as it is visible in the data. Note, we have not propagated systematic uncertainties, so the significance will be less when these are properly propagate. Nevertheless, low and high redshift CC data currently have a preference for different $\Lambda$CDM fitting parameters. Interestingly, as we have shown in section 4, resorting exclusively to MCMC marginalisation one cannot arrive at this conclusion because 2D posteriors are unconstrained, and as a result, uninformative. It is important that observables at all redshifts return similar $H_{0}$ values because if the CC program is claiming an 8% constraint on the Hubble constant, $H_{0}=66.7\pm 5.5$ km/s/Mpc Moresco:2023zys , it is imperative that all subsets of the data are consistent with this result. If they are not, then we are staring at either systematics or model breakdown. Admittedly, demanding self-consistency of subsets of a data set confronted to a model is a high bar, but it is important that data sets result in overlapping constraints on $\Lambda$CDM parameters, otherwise this makes cosmological inferences moot. Note, the $\Lambda$CDM model is largely only well tested in the DE dominated regime $z\lesssim 1$ and at very high redshifts $z\sim 1100$, which leaves a wide expanse of redshifts to be explored in order to confirm or refute the model. Given the existing $\Lambda$CDM tensions Perivolaropoulos:2021jda ; Abdalla:2022yfr , and the hints of evolution in $H_{0}$, $\Omega_{m}$ and $S_{8}$ across assorted probes in the late Universe $z\lesssim 5$, it would be surprising if all discrepancies could be explained away by systematics. [^13] We are open to the possibility, we just consider it a bad bet at the moment. The odds can of course change as observations improve.
@@
As an aside, it is intriguing that CC data has a preference for larger $H_{0}$ best fit values and smaller $\Omega_{m}$ best fit values beyond $z_{\textrm{min}}=0.7$, as this is traditionally the transition redshift between decelerated and accelerated expansion. Moreover, at higher redshifts $z\sim 2.3$, there is not only a longstanding anomaly in Lyman-$\alpha$ BAO duMasdesBourboux:2020pck , but QSOs also show a preference for a lower luminosity distance, $D_{L}(z)$, relative to Planck-$\Lambda$CDM Risaliti:2015zla ; Risaliti:2018reu . Translated into $\Lambda$CDM parameters, this corresponds to conversely larger $\Omega_{m}$ values, e. g. Yang:2019vgk ; Khadka:2020vlh ; Khadka:2020tlm ; Khadka:2021xcc ; Pourojaghi:2022zrh , and consequently smaller $H_{0}$ values. Thus, the emerging probes CC and QSOs Moresco:2022phi do not appear to be in sync on high redshift $\Lambda$CDM inferences. Nevertheless, neither may be inconsistent with the anomaly in Lyman-$\alpha$ BAO. Relative to Planck-$\Lambda$CDM, Lyman-$\alpha$ BAO prefers smaller values of $D_{M}(z):=c\int_{0}^{z}1/H(z^{\prime})\,\textrm{d}z$ and smaller values of $H(z)$ (larger values of $D_{H}(z):=c/H(z)$). [^14] In this statement we assumed the Planck value $r_{d}\sim 147$ Mpc Planck:2018vyg If we reinstate the radius of the sound horizon in these expressions, one recognises that changing the sound horizon, as advocated by early Universe resolutions to Hubble tension, cannot consistently address the Lyman-$\alpha$ BAO anomaly. In general, even for the Planck-$\Lambda$CDM sound horizon, one cannot get both a smaller $D_{M}(z)$ and smaller $H(z)$ from a strictly increasing function, such as the $\Lambda$CDM $H(z)$. As a result, deviations from the Planck-$\Lambda$CDM model that address this anomaly are expected to lead to wiggles in $H(z)$ Akarsu:2022lhx , which are unsurprisingly seen in data reconstructions Zhao:2017cud ; Wang:2018fng ; Escamilla:2021uoj . Finally, evolution in $H_{0},\Omega_{m}$ discussed here cannot be explained or accommodated by early resolutions to Hubble tension relying on a change in the $r_{d}$ at very high $z$.. If CC data prefer less evolution in $H(z)$ in the matter-dominated regime, then this is consistent with the preference for a smaller $H(z)$ from Lyman-$\alpha$ BAO. Furthermore, QSO data prefers smaller luminosity distances $D_{L}(z)$ relative to Planck, which are consistent with the smaller $D_{M}(z)\propto D_{L}(z)$ values preferred by Lyman-$\alpha$ BAO. Thus, even if CC and QSOs appear to be showing diverging behaviour in the cosmological fitting parameters $(H_{0},\Omega_{m})$, this may still turn out to be consistent with Lyman-$\alpha$ BAO. We await future DESI DESI:2023ytc data releases to ascertain if the non-evolving $H(z)$ trend in high redshift CC data is physical or not.
@@
Finally, we come to our third and main result outlined in section 5. We have revisited a $\sim 2\sigma$ tension between high redshift CC and BAO data reported in Colgain:2022rxy , where the significance was estimated through mock simulations. Here, we have upgraded the BAO data to the latest constraints and again recover a $>2\sigma$ discrepancy in $(H_{0},\Omega_{m})$ with different methodology. This provides a consistency check that there is evolution in cosmological fitting parameters from OHD between low and high redshifts in the late Universe. Note, this evolution runs contrary to the non-evolving $H(z)$ seen in high redshift CC data because it assumes Planck has accurately constrained the high redshift behaviour of the Hubble parameter in (1). Nevertheless, both with and without a Planck prior on $\Omega_{m}h^{2}$, evolution at $\gtrsim 2\sigma$ is evident in OHD data. It should be stressed that evolution is evident in PDFs of best fit $\Lambda$CDM parameters fitted to a large number of Planck-$\Lambda$CDM mocks Colgain:2022tql , so evolution in fitting parameters from observed data can be expected. It is imperative to revisit the remaining observations in Colgain:2022rxy ; Malekjani:2023dky in order to confirm the significance of $\sim 2\sigma$ hints of evolution found separately in Type Ia SNe and QSO data sets.